Find the length and width of a rectangle lot with a perimeter of 70 meters if the length is 5 times more than the width.
Answer by Mathtut(3670)

let L and W be the length and width
:
P=2L+2W......eq 1
L=5W.........eq 2
P=70.........eq 3
:
Plug P's value from eq 3 and L's value from eq 2 into eq 1
:
70=2(5W)+2W
:
70=10W+2W
:
70=12W
:
W=70/12=35/6 m
:
L=5(35/6)=175/6 m